OpenStudy (anonymous):

Use the square root property to answer (x+8)^2=81

OpenStudy (anonymous):

I x+8 I =sqrt(81) x+8=9 x=9-8=1 or I x+8 I =-sqrt(81) x+8=-9 x=-9-8=-17

OpenStudy (anonymous):

right